OpenStudy (ivancsc1996):

Can somebody help me understand this equation? A graph of y(x) would be very helpful. It arrised from solving newton's equations for a uniform gravitational field under the presence of a resiatance force proportional to the particle's velocity

OpenStudy (ivancsc1996):

\[y(x)=(\frac{ gm }{ bv _{0} }\sec \theta+\tan \theta)x+\frac{ gm ^{2} }{ b ^{2} }\ln (\frac{ m }{ b }v _{0} \cos \theta -x)\]